a. Find the volume V(S) of the solid S by revolving the region R bounded by x 2 + y − 4 = 0 and x − y + 2 = 0 about y = 0. (6 pts.) b. Set-up the integral that represents V(S) when the region R in (a.) is revolved about x = −2. (4 pts.)


1
Expert's answer
2022-04-11T07:29:06-0400

a)

Points of interception:

"x^2+y-4=0\\x-y+2=0"

"x^2+x+2-4=0"

"x=-2" , "x=1" .



V(S) can be obtained if we subtract volume

bounded by "x-y+2=0" , "y=0" , "-2\le x\le1" and revolved about "y=0"

from the volume

bounded by "x^2+y-4=0" , "y=0" , "-2\le x\le1" and revolved about "y=0"

"V(S)=\pi\int_{-2}^1((-x^2+4)^2-(x+2)^2)dx" "=\pi\int_{-2}^1(x^4-8x^2+16-x^2-4x-4)dx"

"=\pi(x^5/5-9x^3/3-4x^2/2+12x)|_{-2}^1" "=\pi(1/5-3-2+12+32/5-3\cdot8""+2\cdot4+12\cdot2) =108\pi/5"

b)

Let’s move to new coordinates:

"t=x+2"

u=y

Now region R is bounded by

"(t-2)^2+u-4=0" and "t-u=0"

V(S) is obtained by revolving R about "t=0"



Look at the picture. Region R is divided into R1 and R2.

Volume V1 is obtained by revolving R1 about "t=0" can be calculated by subtracting volume

bounded by "t=2-\sqrt{4-u}" , "t=0" , "0\le u\le3" and revolved about "t=0"

from the volume

bounded by "t=u" , "t=0" , "0\le u\le3" and revolved about "t=0".

"V1=\int_0^3\pi(u^2-(2-\sqrt{4-u})^2)du"

Volume V2 is obtained by revolving R2 about "t=0" can be calculated by subtracting volume bounded by "t=2-\sqrt{4-u}" , "t=0" , "3\le u\le4" and revolved about "t=0"

from the volume

bounded by "t=2+\sqrt{4-u}" , "t=0" , "3\le u\le4" and revolved about "t=0".

"V2=\int_3^4\pi((2+\sqrt{4-u})^2-(2-\sqrt{4-u})^2)du"

"V(S)=V1+V2=""\int_0^3\pi(u^2-(2-\sqrt{4-u})^2)du+""\int_3^4\pi((2+\sqrt{4-u})^2-(2-\sqrt{4-u})^2)du"